A 2m long fixed string has nodes at both ends when it vibrates in its fundamental
mode. If the distance between this nodes is 12cm. what are the number of (1) nodes
(b) anti-nodes?
1
Expert's answer
2017-12-07T10:12:07-0500
a) The number of nodes is
N+1=2/(0.1 )+1=21
b) The number of antinodes is
N=2/(0.1 )=20